# Missing Data Imputation by Reducing Mutual Information with Rectified Flows

### v2

@@ 5.3 Contemporary Approaches: Learn Jointly, Sample/Inpaint Conditionally
Traditionally, learning a high-dimensional joint distribution $q(\mathbf{x})$ is intractable due to its normalising constant. Recently, it was realized that one can learn a joint model $q(\mathbf{x})$ without dealing with the normalising constant. (Uehara et al., 2020) first estimates the joint model $q$ via Score Matching (Hyvärinen, 2005) or Noise Contrastive Estimation (Gutmann and Hyvärinen, 2010) , then sample from $q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$ via importance sampling. This method works well when $\mathbf{x}_{1-\mathbf{m}}$ is in a relatively low-dimensional space. (Chen et al., 2024) noticed that given a joint model $q$, this conditional sampling problem could be solved using Stein Variational Gradient Descent (SVGD) (Liu and Wang, 2016). However, SVGD utilizes properties of RKHS function, and require kernel computations. Therefore, does not scale to high-dimensional datasets.
In recent years, it is also realized that one can “inpaint” samples given a diffusion model trained on the joint samples (Lugmayr et al., 2022). This idea gives rise to DiffPuter (Zhang et al., 2025), an imputer that trains a joint diffusion model to sample from $q$. Then use an inpainter to impute missing values given the observed vector $\mathbf{x}_{\mathbf{m}}$. However, an inpainter is not a standard backward process of diffusion dynamics, thus in general, is not guaranteed to generate samples from $q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$.
@@ 6 Experiments
Details on experimental settings and implementations can be found in Appendix E. Additional results can be found in Appendix F. In experiments, we match model architectures across all methods, if possible. For example, if our method uses CNN, we use the same architectures for other neural-network-based methods as well.
@@
Figure 1: Pairwise density plots of the imputed data and the true data and MMD/MI of MIRI.
@@ 6.1 Toy Example: Assessing Imputation Quality Using MMD and Mutual Information
GAIN (Yoon et al., 2018) 164.11 21.21 0.7803 281.62 16.20 0.5576 285.53 11.99 0.2933
KnewImp (Chen et al., 2024) 153.09 18.84 0.6463 193.68 15.81 0.4740 264.40 14.04 0.3317
MissDiff (Ouyang et al., 2023) 90.51 22.29 0.7702 129.84 19.65 0.6648 197.91 16.78 0.4989
HyperImpute (Jarrett et al., 2022) 8.92 34.09 0.9750 65.01 23.22 0.7931 130.36 20.17 0.6533
@@
In this section, we showcase the performance of the proposed MIRI imputer on a toy example that is a mixture of 2-dimensional Gaussian: $\mathbf{X}^{*}\sim\mathcal{N}([-2,-2],0.5^{2}\mathbf{I})+\mathcal{N}([2,2],0.5^{2}\mathbf{I})$. The missing mask is generated as $\mathbf{M}\sim\prod_{j=1}^{d}\text{Bernoulli}_{M_{j}}(0.7)$, i.e., 30% of the data is missing completely at random. We draw 6000 i.i.d. samples from $\mathbb{P}_{\mathbf{X}^{*}}$ and $\mathbb{P}_{\mathbf{M}}$ to create missing observations. The imputed samples are visualized in the pairwise density plots in Figure 1. We choose HyperImpute (Jarrett et al., 2022) and KnewImp (Chen et al., 2024) as comparisons. It can be seen that MIRI not only captures the bimodal structure of the true data but also preserves the proportion of the two modes. On the other hand, HyperImpute and KnewImp fail to recover the correct proportion of the two modes. We also show performance metrics (MAE, RMSE, MMD) with various other imputation algorithms (such as MissDiff, GAIN, TabCSDI (Zheng and Charoenphakdee, 2022)) in Table 3 in the Appendix F.1. Finally, in Figure 1(d), we show that both MIRI’s MMD and MI decrease over iterations, proving that MI between the imputed data and the missing mask is a good criterion for assessing how well the imputed data recovers $\mathbb{P}_{\mathbf{X}^{*}}$.
6.2 Real-World Tabular Imputation: UCI Regression Benchmarks
We evaluate MIRI on 10 standard UCI regression benchmarks (Kelly et al., 2017). We select these datasets for their diversity in sample size and dimensionality (see Table 2 in Appendix E.1). We simulate missingness by independently masking 20%, 40%, and 60% of entries under an MCAR mechanism. Our method is compared with HyperImpute, KnewImp and TabCSDI. Figure 2 reports the average MMD over 10 trials (error bars denote one standard deviation). As shown, even in the 60% missing data cases, MIRI outperforms or performs comparably to all competitors. Results for the 20% and 40% missingness scenarios, as well as additional comparisons, are provided in Appendix F.2.
@@ 6.3 Real-World Image Imputation: CIFAR-10 and CelebA Benchmarks
(a) 15 uncurated 32$\times$32 CIFAR-10 images and their imputations. Pixels are removed from all RGB channels.
(b) 15 uncurated 64$\times$64 CelebA images and their imputations. Pixels are removed from each RGB channel independently.
@@
We first assess MIRI on the high-dimensional CIFAR-10 dataset (Krizhevsky et al., 2009) (32$\times$32 RGB images) by randomly removing 60% of pixels (all RGB channels at once), and training on just 5 000 samples (< 10% of the full set). Figure 3(a) shows the imputation reconstructions at 60% missingness of the generative model-based approaches GAIN, KnewImp, and MissDiff, alongside our method. Table 1 reports FID, PSNR and SSIM at 20%, 40% and 60% missingness. Across this range, MIRI achieves 65.4 – 93.4% lower FID, 38.1 – 44.9% higher PSNR and 26.4 – 61.6% higher SSIM than the strongest flow-based alternative. These margins remain significant even at 60% missingness, demonstrating robustness to severe sparsity. We have also included results for HyperImpute in Table 1, but in the main text, we only show visual reconstructions for generative model-based approaches. Visual reconstructions for HyperImpute, along with additional results, are provided in Appendix F.3.
@@
Additionally, we evaluate the performance of MIRI on a higher-resolution dataset on CelebA (Liu et al., 2015) (64$\times$64 RGB images), where we adopt a complementary missingness mechanism that independently masks each RGB channel. The imputation reconstruction of images using 5 000 samples displayed in Figure 3(b) substantiates MIRI’s strong performance.
7 Limitations and Future Work
@@
Our method is computationally heavy, as for each MIRI iteration, we need to run a full rectified flow algorithm, including training the velocity field and solving the ODE. One interesting future work is to incorporate recent distillation methods to accelerate sample generation.
@@
Our Theorem 1 is a population-level statement, meaning it assumes the population optimal $\mathbf{v}^{*}$ is obtained from the training objective (6) and that the imputation ODE (7) is solved exactly. It does not consider errors due to finite sample approximation, optimisation and inaccurate ODE solver. How to extend our theoretical results to the finite-sample optimal $\hat{\mathbf{v}}$ and the approximate imputation ODE solver (like the Euler method) is another avenue for future works.
Acknowledgements
Funding for this work was provided by the University of Bristol Undergraduate Summer Research Bursary 2024. Jiahao Yu was funded by the Alumni Foundation, and Leyang Wang by the Heilbronn Institute for Mathematical Research. Partial computational resources were provided by the Advanced Computing Research Centre, University of Bristol (http://www.bristol.ac.uk/acrc/).
Appendix A Proof of Proposition 1
@@
Consider the difference between the mutual information at iteration $t$ and $t-1$:
$\mathrm{D}\left[\mathbb{P}_{\mathbf{X},\mathbf{M}}^{(t)}|\mathbb{P}_{\mathbf{X}}^{(t)}\otimes\mathbb{P}_{\mathbf{M}}\right]-\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t-1)},\mathbf{M}}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$
$=$ $\underbrace{\mathrm{D}\left[\mathbb{P}_{\mathbf{X},\mathbf{M}}^{(t)}|\mathbb{P}_{\mathbf{X}}^{(t)}\otimes\mathbb{P}_{\mathbf{M}}\right]-\mathrm{D}\left[\mathbb{P}_{\mathbf{X},\mathbf{M}}^{(t)}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]}_{A}+\underbrace{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{D}\left[\mathbb{P}_{\mathbf{X},\mathbf{M}}^{(t)}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]}-{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t-1)},\mathbf{M}}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]}}_{B}.$ (11)
Since $\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$ is minimized at $\mathbf{g}_{t}$ and $\mathbf{X}^{(t)}=\mathbf{X}^{(t)}(\mathbf{g}_{t})$ if $\mathbf{g}_{t}$ is optimal,
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{D}\left[\mathbb{P}_{\mathbf{X},\mathbf{M}}^{(t)}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]}\leq\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right],\forall\mathbf{g}.$
Moreover, ${\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t-1)},\mathbf{M}}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]}=\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]|_{\mathbf{g}=\mathrm{Identity}}$, so $B$ is non-positive. To show $A$ is non-positive, we write out the KL-divergences in $A$. Let $p(\mathbf{x},\mathbf{m})$ be the density of $(\mathbf{X}^{(t)},\mathbf{M})$ and $q(\mathbf{x})$ be the density of $\mathbf{X}^{(t-1)}$.
$\mathrm{D}\left[\mathbb{P}_{\mathbf{X},\mathbf{M}}^{(t)}|\mathbb{P}_{\mathbf{X}}^{(t)}\otimes\mathbb{P}_{\mathbf{M}}\right]-\mathrm{D}\left[\mathbb{P}_{\mathbf{X},\mathbf{M}}^{(t)}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$
$=$ $\mathbb{E}\left[\log\frac{p(\mathbf{X}^{(t)},\mathbf{M})}{p(\mathbf{X}^{(t)})p(\mathbf{M})}\right]-\mathbb{E}\left[\log\frac{p(\mathbf{X}^{(t)},\mathbf{M})}{q(\mathbf{X}^{(t)})p(\mathbf{M})}\right]$
$=$ $-\mathbb{E}\left[\log p(\mathbf{X}^{(t)})\right]+\mathbb{E}\left[\log q(\mathbf{X}^{(t)})\right]\leq 0,$
@@
Appendix B Proof of Theorem 1

### v3

@@ 4.3 Contemporary Approaches: Learn Jointly, Sample/Inpaint Conditionally
Traditionally, learning a high-dimensional joint distribution $q(\mathbf{x})$ is intractable due to its normalising constant. Recently, it was realized that one can learn a joint model $q(\mathbf{x})$ without dealing with the normalising constant. (uehara20b) first estimates the joint model $q$ via Score Matching (Hyvaerinen2005) or Noise Contrastive Estimation (gutmann10a) , then sample from $q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$ via importance sampling. This method works well when $\mathbf{x}_{1-\mathbf{m}}$ is in a relatively low-dimensional space. (chen2024rethinking) noticed that given a joint model $q$, this conditional sampling problem could be solved using Stein Variational Gradient Descent (SVGD) (LiuQ2016SVGD). However, SVGD utilizes properties of RKHS function, and require kernel computations. Therefore, does not scale to high-dimensional datasets.
In recent years, it is also realized that one can “inpaint” samples given a diffusion model trained on the joint samples (Lugmayr_2022_CVPR). This idea gives rise to DiffPuter (zhang2025diffputer), an imputer that trains a joint diffusion model to sample from $q$. Then use an inpainter to impute missing values given the observed vector $\mathbf{x}_{\mathbf{m}}$. However, an inpainter is not a standard backward process of diffusion dynamics, thus in general, it is not guaranteed to generate samples from $q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$.
@@ 5 Experiments
This section evaluates MIRI’s performance on a variety of imputation tasks. Details on experimental settings and implementations can be found in Appendix G. Additional results can be found in Appendix H. To ensure fair comparisons, we match model architectures across all methods wherever possible. For example, if our method uses CNN, we use the same architectures for other neural-network-based methods as well.
Imputation methods aim to either predict the conditional mean $\mathbb{E}[\mathbf{X}_{1-\mathbf{m}}|\mathbf{X}_{\mathbf{m}}]$, evaluated by metrics like MAE/RMSE, or learn to sample from the conditional distribution $\mathbb{P}_{\mathbf{X}_{1-\mathbf{m}}|\mathbf{X}_{\mathbf{m}}}$, evaluated by the distributional fidelity. Our method, MIRI, is designed for the latter. Thus, we mostly evaluate the performance by using metrics that measure the distributional fidelity (such as MMD (gretton2012kernel)).
@@
Figure 1: Pairwise density plots of the imputed data and MMD/MI of MIRI.
@@ 5.1 Toy Example: Assessing Imputation Quality Using MMD and Mutual Information
In this section, we showcase the performance of the proposed MIRI imputer on a toy example that is a mixture of 2-dimensional Gaussian: $\mathbf{X}^{*}\sim 0.5\mathcal{N}([-2,-2],0.5^{2}\mathbf{I})+0.5\mathcal{N}([2,2],0.5^{2}\mathbf{I})$. The missingness mask is generated as $\mathbf{M}\sim\prod_{j=1}^{d}\text{Bernoulli}_{M_{j}}(0.7)$, i.e., 30% of the data is Missing Completely at Random. We draw 6000 i.i.d. samples from $\mathbb{P}_{\mathbf{X}^{*}}$ and $\mathbb{P}_{\mathbf{M}}$ to create missing observations. The imputed samples are visualized in the pairwise density plots in Figure 1. We choose HyperImpute (Jarrett2022HyperImpute) and KnewImp (chen2024rethinking) as comparisons. It can be seen that MIRI not only captures the bimodal structure of the true data but also preserves the proportion of the two modes. On the other hand, HyperImpute and KnewImp fail to recover the correct proportion of the two modes. We also show performance metrics (MAE, RMSE, MMD) with various other imputation algorithms (such as MissDiff, GAIN, TabCSDI (zheng_diffusion_2023)) in Table 5 in the Appendix H.1. Finally, in Figure 1(d), we show that both MIRI’s MMD and MI decrease over iterations, proving that MI between the imputed data and the missingness mask is a good criterion for assessing how well the imputed data recovers $\mathbb{P}_{\mathbf{X}^{*}}$.
5.2 Real-World Tabular Imputation: UCI Regression Benchmarks
Table 1: Aggregated MMD rankings for UCI (40% missingness) across 10 datasets. Lower is better.
Method MCAR MAR MNAR
TabCSDI (zheng_diffusion_2023) 6.4 6.4 6.1
GAIN (yoon2018gain) 6.2 6.2 5.9
TDM zhao2023transformed 4.5 4.4 4.2
KnewImp (chen2024rethinking) 4.1 4.2 3.8
MIWAE (mattei2019miwae) 3.0 2.9 2.8
HyperImpute (Jarrett2022HyperImpute) 1.6 1.5 1.6
MIRI (Ours*) 1.4 2.0 1.3
@@ 5.2 Real-World Tabular Imputation: UCI Regression Benchmarks
GAIN (yoon2018gain) 164.11 21.21 0.7803 281.62 16.20 0.5576 285.53 11.99 0.2933
KnewImp (chen2024rethinking) 153.09 18.84 0.6463 193.68 15.81 0.4740 264.40 14.04 0.3317
MissDiff (ouyang_missdiff_2023) 90.51 22.29 0.7702 129.84 19.65 0.6648 197.91 16.78 0.4989
HyperImpute (Jarrett2022HyperImpute) 8.92 34.09 0.9750 65.01 23.22 0.7931 130.36 20.17 0.6533
@@
We evaluate MIRI on 10 standard UCI regression benchmarks (kelly2019). We select these datasets for their diversity in sample size and dimensionality (see Table 4 in Appendix G.1). Following (muzellec2020missing), we simulate missingness under MCAR/MAR/MNAR mechanisms. We simulate 20%, 40%, and 60% missingness for MCAR and MNAR, and 40% and 80% for MAR. For the MAR mechanism, the missingness mask is conditioned on a randomly selected 50% of the variables. In the main paper, we compare our method with HyperImpute, KnewImp, TabCSDI, GAIN, MIWAE (mattei2019miwae), and TDM (zhao2023transformed).
Figure 2 reports the average MMD over 10 trials (error bars denote one standard deviation). As shown, even in the 60% missing data cases, MIRI outperforms or performs comparably to all competitors. We aggregate performance using mean rank across 10 datasets. As shown in Table 1 for the 40% missingness setting, MIRI attains the best mean rank under MCAR and MNAR settings and is highly competitive under MAR setting.
Comprehensive results, including performance across all specified missingness levels and comparisons with an extended set of baseline methods, are deferred to Appendix H.2.
@@ 5.3 Real-World Image Imputation: CIFAR-10 and CelebA Benchmarks
(a) 15 uncurated $\times32$ CIFAR-10 images and their imputations. Pixels are removed from all RGB channels.
(b) 15 uncurated $\times64$ CelebA images and their imputations. Pixels are removed from each RGB channel independently.
@@
Table 3: Accuracy of 10-class CIFAR-10 classification on imputed data at varying missingness. Higher is better.
Method 20% 40% 60%
KnewImp (chen2024rethinking) 0.267 0.137 0.106
GAIN (yoon2018gain) 0.337 0.133 0.096
MissDiff (ouyang_missdiff_2023) 0.486 0.269 0.152
HyperImpute (Jarrett2022HyperImpute) 0.804 0.405 0.212
MIRI (Ours*) 0.812 0.525 0.364
We assess MIRI on high-dimensional image datasets.
First, on CIFAR-10 (krizhevsky_cifar-10_2009) ($\times32$ RGB), we train using only 5 000 samples (<10% of the full set) with 60% of pixels randomly removed (all RGB channels jointly). Figure 3(a) compares MIRI’s 60% missingness reconstructions against generative approaches (GAIN, KnewImp, MissDiff). Table 2 reports FID, PSNR, and SSIM at 20%, 40%, and 60% missingness, showing MIRI strongly outperforms the best generative alternatives across this range, even in challenging cases.
See Appendix H.3 for additional results, and Table 3 for downstream classification results.
@@
Second, we test MIRI on higher-resolution CelebA (liu2015faceattributes) ($\times64$ RGB) with 5 000 samples, using a mechanism that masks each RGB channel independently. Figure 3(b) again substantiates MIRI’s strong imputation performance.
6 Limitations and Future Works
Despite the competitive results, MIRI has several limitations that open avenues for future work.
@@ 6 Limitations and Future Works
A primary limitation is computational intensity. Each MIRI iteration requires running a full rectified flow algorithm, including training a velocity field and solving the corresponding ODE. One interesting future work is to incorporate recent distillation methods to accelerate sample generation.
@@
Our Theorem 1 is a population-level result, meaning it assumes the optimal velocity field $\mathbf{v}^{*}$ is obtained and the imputation ODE is solved exactly. It does not consider errors due to finite sample approximation, optimisation and inaccurate ODE solver. Extending our theoretical results to the finite-sample optimal $\hat{\mathbf{v}}$ and the approximate imputation ODE solver (like the Euler method) is another avenue for future works.
A more fundamental limitation is that our objective, which minimizes the mutual information between the data and the mask, is not designed for Missing Not at Random (MNAR) scenarios where the missingness mechanism depends on the true values. This is a common challenge for many modern imputation frameworks, and extending our method to handle such cases is a crucial direction.
Finally, MIRI could be extended to a self-supervised framework for time-series. By training on artificially masked, fully-observed sequences, we could learn a general-purpose imputation vector field. The key advantage here is flexibility: because MIRI’s vector fields are agnostic to missingness patterns, this approach could yield a single, pre-trained foundation model for robust, general-purpose time-series imputation and forecasting.
7 Conclusion
In this work, we introduced Mutual Information Reducing Iterations (MIRI), a framework that reduces data-mask dependency by iteratively minimizing $\mathrm{D}[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}]$. We proved that MIRI decreases mutual information $\mathrm{I}(\mathbf{X};\mathbf{M})$. The optimal imputer is constructed by solving an ODE trained with a rectified flow objective. We also showed that existing imputation methods such as GAIN are approximate special cases of this approach. Empirically, MIRI also demonstrated strong distributional fidelity across synthetic, tabular, and image benchmarks.
Acknowledgments
We thank the four anonymous reviewers and the area chair for their insightful comments and suggestions. SL thanks Josh Givens, Sam Power, Katarzyna Reluga and other colleagues at Bristol Mathematics for helpful discussions. Partial funding for this work was provided by the University of Bristol School of Mathematics Summer Research Bursary 2024; JY was funded by the Alumni Foundation, and LW by the Heilbronn Institute for Mathematical Research. JY also acknowledges support from the Scholar Award and the ED Davies Travel Fund, provided by the Neural Information Processing Foundation and the Fitzwilliam College, University of Cambridge, respectively. Partial computational resources were provided by the Advanced Computing Research Centre at the University of Bristol (https://www.bristol.ac.uk/acrc/).
Appendix A Illustration of Sequential Imputation
Figure 4: Schematic of the sequential imputation algorithm. The process begins with an initial imputation, $\mathbf{X}^{(0)}$, and iteratively refines the missing values over $T$ steps.
Figure 4 illustrates the sequential imputation framework, where a model is iteratively trained to refine estimates for missing data.
Appendix B Proof of Proposition 1
@@
Consider the difference between the mutual information at iteration $t$ and t1t-1:
$\mathrm{D}\left[\mathbb{P}_{\mathbf{X},\mathbf{M}}^{(t)}\|\mathbb{P}_{\mathbf{X}}^{(t)}\otimes\mathbb{P}_{\mathbf{M}}\right]-\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t-1)},\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$
$=$ $\underbrace{\mathrm{D}\left[\mathbb{P}_{\mathbf{X},\mathbf{M}}^{(t)}\|\mathbb{P}_{\mathbf{X}}^{(t)}\otimes\mathbb{P}_{\mathbf{M}}\right]-\mathrm{D}\left[\mathbb{P}_{\mathbf{X},\mathbf{M}}^{(t)}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]}_{A}+\underbrace{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{D}\left[\mathbb{P}_{\mathbf{X},\mathbf{M}}^{(t)}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]}-{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t-1)},\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]}}_{B}.$ (16)
Since $\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$ is minimized at $\mathbf{g}_{t}$ and $\mathbf{X}^{(t)}=\mathbf{X}^{(t)}(\mathbf{g}_{t})$ by definition,
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{D}\left[\mathbb{P}_{\mathbf{X},\mathbf{M}}^{(t)}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]}\leq\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right],\forall\mathbf{g}.$ (17)
Moreover, ${\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t-1)},\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]}=\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]|_{\mathbf{g}=\mathrm{Identity}}$, so $B$ is non-positive. To show $A$ is non-positive, we write out the KL-divergences in $A$. Let $p(\mathbf{x},\mathbf{m})$ be the density of $(\mathbf{X}^{(t)},\mathbf{M})$ and $q(\mathbf{x})$ be the density of $\mathbf{X}^{(t-1)}$.
$\mathrm{D}\left[\mathbb{P}_{\mathbf{X},\mathbf{M}}^{(t)}\|\mathbb{P}_{\mathbf{X}}^{(t)}\otimes\mathbb{P}_{\mathbf{M}}\right]-\mathrm{D}\left[\mathbb{P}_{\mathbf{X},\mathbf{M}}^{(t)}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$ (18)
$=$ $\mathbb{E}\left[\log\frac{p(\mathbf{X}^{(t)},\mathbf{M})}{p(\mathbf{X}^{(t)})p(\mathbf{M})}\right]-\mathbb{E}\left[\log\frac{p(\mathbf{X}^{(t)},\mathbf{M})}{q(\mathbf{X}^{(t)})p(\mathbf{M})}\right]$ (19)
$=$ $-\mathbb{E}\left[\log p(\mathbf{X}^{(t)})\right]+\mathbb{E}\left[\log q(\mathbf{X}^{(t)})\right]\leq 0,$ (20)
@@
Appendix C Proof of Theorem 1